The perimeter of a rectangle is 38 centimeters. The length is 7.5 centimeters. What is the width?
Answer by Aust05(12)

the formula for perimeter is P= 2L + 2W so L= 7.5 so 2* 7.5 = 15 then subtract that from 38 getting 13. so 2W=13 divide both sides by 2 to find that W=6.5 so 2W+2L=38 2w+ 2*7.5=38 2w + 15= 38 then 2w=38-15 then 2W=13 w=13/2 w=6.5